OpenStudy (anonymous):

a forklift lifting a crate of mass 100kg at a constant velocity to a height of 4 m over a time of 4s. The forklift then holds in place for 20s. How much power does the forklift exert in lifting the crate.

OpenStudy (anonymous):

p=work/time; w=force X distance. w=100(9.8) X 4= 3920 p=3920/4=980, dont count the 20 seconds because no work is being done.